Both cordless drills are missing from the tool chest and no one has signed them out in the logbook. Please return right away – someone else needs to use one.
All equipment status updates.
Maintenance is complete. The instrument is working great.
Okay to use it!
The TEM is ready to be used.
We got very lucky and saved the ion pump. The flake moved when I vented the column.
I suggest you use it soon before any other problems develop.
Just a reminder that Hitachi will be in town to service the FESEM beginning Monday afternoon. He should finish by Thursday.
We switched over to the new logbook software this morning in the M&M labs. ATDC and Chem Sci haven’t been done yet but will before week is done.
Several important things to note.
1. your accounts won’t be associated with your ID yet. We will have to do that as we go. The moral is – don’t wait until the last minute to do this in case no one is nearby to help you. This is especially the case if you come after hours.
2. there is one database now instead of 9. If we put your name in one computer it is valid for them all. we will assign scopes to you based on your training.
3. when you login you MUST chose the correct microscope. I’m going to ask if we can have the system default to your “scope(s)”, but for now pay attention to the scope you are trying to login to.
Danielle, Liz and I are ready to help.
The FESEM software had a crude tool for making on-screen measurements, but there is a software program, PCI, that can be used for improved results. If you are interested in knowing more about this program let me know. It is quite easy to use.
I’m going to take vacation on Fridays again the summer beginning May 10. Danielle and Liz will be here part of that day and I’ll normally be in town so I could come back to work in case of emergency.
I have put a band around the main valve door to keep the lever closed. It is easy to push it out of the way when you want to open the door. Please put it back when you quit your session.
The manufacturer is coming to service the FESEM next week. The work will begin Monday night and the microscope will be down until Thursday or Friday. The FIB will also be serviced while he is here but that will not affect most of you.
3:35 p.m. update. I adjusted a switch position and it may be okay now. If you hear a rattle noise push the MV1 lever (airlock valve) closed tightly and let me know right away.
2:42 p.m. The pneumatic lock on the FESEM that holds the Exchange Door closed is not functioning properly. The chamber vacuum will drop when the door drifts open and may cause the beam to automatically shut off. Please keep an eye on the door to make sure it remains in the closed position (EVAC Light will be a solid green and SC Vacuum Light will be green), especially while imaging.
We are contacting Hitachi to get a replacement.